Direct Labor Variances

La Batre Bicycle Company manufactures commuter bicycles from recycled materials. The following data for July of the current year are available:

Quantity of direct labor used 470 hrs.
Actual rate for direct labor $10.6 per hr.
Bicycles completed in April 220 bicycles
Standard direct labor per bicycle 2 hrs.
Standard rate for direct labor $10.8 per hr.

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance $ Favorable
Direct Labor Time Variance $ Unfavorable
Total Direct Labor Cost Variance $ Unfavorable

b. How much direct labor should be debited to Work in Process?
$
Answer #1

Labor rate variance = (SR-AR) * AH

= (10.8-10.6)*470

= - 94 Favourable (minus)

Labor time variance = (SH-AH) * SR

= (220*2 - 470) * 10.8

= 324 Unfavorable

Total labor variance

= 94F+324 U

= 230 Unfavorable

B

Direct labor debited to wip

= 220*2*10.8

= 4752